OpenStudy (anonymous):

I made an equation and I know the equation will work for it but I need to find out what x equals. 500=[2pi(x/2)]2+x

OpenStudy (unklerhaukus):

\[500=\big[2\pi(x/2)\big]^2+x\]

OpenStudy (anonymous):

Oops my bad hold on i'll write it correctly.

OpenStudy (anonymous):

\[500=[2\pi(x/2)]2+x\]

OpenStudy (anonymous):

so it's times 2 not to the power of 2

OpenStudy (unklerhaukus):

oh \[500=\big[2\pi(x/2)\big]\times2+x\]

OpenStudy (anonymous):

not sure what to do after I divide both sides by 2

OpenStudy (anonymous):

Yup

OpenStudy (unklerhaukus):

lets look at this bit first \[[2π(x/2)]× 2\], can you simplify this

OpenStudy (anonymous):

not sure what I did was I divided both sides by 2 to get rid of the 2 but I don't think that was right

OpenStudy (anonymous):

\[4\pi(x/2)\]

OpenStudy (unklerhaukus):

Yes, \[[2π(x/2)]×2=4\pi (x/2)\] now, simplify again \[=\]

OpenStudy (unklerhaukus):

(combine the 4 and the /2 )

OpenStudy (anonymous):

\[2pix\]

OpenStudy (unklerhaukus):

yeah so you have \[500=\big[2\pi(x/2)\big]\times2+x\\ 500= 2\pi x+x\]

OpenStudy (unklerhaukus):

now can you see a common factor in the right hand side ?

OpenStudy (anonymous):

Oh ok that makes sense I got ti from here thanks a bunch!